OpenStudy (anonymous):

A motor requires 400. joules of energy to lift a 5.0 kg mass 2.0 meters. Calculate the efficiency of this motor

OpenStudy (quinton1011):

what do you think the answer is?

OpenStudy (anonymous):

well im sure its not 25% @quinton1011

OpenStudy (quinton1011):

okay

OpenStudy (quinton1011):

is there any answer choices?

OpenStudy (quinton1011):

hello?

OpenStudy (quinton1011):

energy input = 400 J energy output = 98 J (from problem 2) efficiency = (energy output / energy input) * 100 efficiency = (98 / 40) * 100 = 24.5% medal if helped

OpenStudy (anonymous):

sorry the answers were 25%, 12% and 10%. So you got 25%? thanks

OpenStudy (raffle_snaffle):

@Quinton1011 don't just give the answer away...

OpenStudy (anonymous):

from where u taken o/p is 98???

OpenStudy (anonymous):

The answer is 25%